The length of the bases of a trapezoid are 18 centimeters and 30 centimeters. Find the length of its median.

1
Expert's answer
2022-03-14T10:50:23-0400

The length of its median is equal to

"m=\frac{a+b}2=\frac{18+30}2=24."